1/(3x +5x^-n) as integration sir please solve it fast
 

Grade:12

1 Answers

Aditya Gupta
5 years ago
1/(3x +5x^-n) = x^n/(3x^(n+1)+5)
now let y=(3x^(n+1)+5)
so dy/dx=3(n+1)x^n
so integral is
dy/3y(n+1)
=lny/3(n+1)
=ln(3x^(n+1)+5)/3(n+1) + C